# On the non-Markovian quantum control dynamics

### v3

@@ III Time-varying linear quantum control dynamics
In the following, we only consider the circumstance that $4\kappa_{n}S_{n}-Q_{n}^{2}\leq 0$, thus the interaction between the atom and environment finally converges to a Markovian form. Return to Eq. (12), we analyze how the quantum dynamics can be influenced by detunings between the atom and cavity, and drive fields applied upon the quantum system.
@@
$=$ $\begin{bmatrix}-L_{n}^{(1)}&-i\tilde{g}_{n}&i\tilde{g}_{n}\\ -i\tilde{g}_{n}&-L_{n}^{(2)}&0\\ i\tilde{g}_{n}&0&-L_{n}^{(3)}\end{bmatrix}$
$+{\rm{diag}}\left(L_{n}^{(1)}-\left(F_{n}+F_{n}^{*}\right),L_{n}^{(2)}-F_{n}^{*},L_{n}^{(3)}-F_{n}\right),$
@@
For the special case where $g_{n}=0$, meaning the multi-level atom only interacts with the non-Markovian environment with $C_{n}=D_{n}^{\rm T}=\textbf{0}_{1\times 3}$, we then construct the vector representing the populations of a multi-level atom as
$\tilde{X}=\left[\langle\sigma^{+}_{1}\sigma^{-}_{1}\rangle,\cdots,\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle,\cdots,\langle\sigma^{+}_{N-1}\sigma^{-}_{N-1}\rangle\right]^{\rm T},$
which is governed by the following real-valued equation if initially $\tilde{X}(0)$ is real,
$\frac{\mathrm{d}}{\mathrm{d}t}\tilde{X}=\left(\textbf{L}+\textbf{P}(t)\right)\tilde{X}(t),$ (36)
@@
then we can derive the following results.
@@
The convergence of quantum dynamics to a Markovian regime means the convergence of $F_{n}(t)$ according to Remark 1. Consequently, the time-varying components in Eq. (36) converge to zero. In this case, $\lim_{t\rightarrow\infty}\textbf{P}(t)=\textbf{0}_{(N-1)\times(N-1)}$. Thus $\Phi\left(t,t_{0}\right)$ is bounded because $\lim_{t\rightarrow\infty}\frac{\partial\Phi\left(t,t_{0}\right)}{\partial t}=0$. ∎
@@
where the non-Markovian interaction between the atom and the environment can influence the integration kernel in Eq. (38), and further influence the convergence rate and steady value of $\tilde{X}$. The dynamics of the equation with the format of Eq. (38) has been introduced in Ref. [58]. Generalized from Theorem 2 in Ref. [58] (Chapter 2, Page 36), we have the following corollary.
@@ III-A $\Delta_{n}\neq 0$, $g_{n}\neq 0$
In the following, we consider the most general case by separating $X(t)$ into its real and imaginary parts. Considering that in Eq. (9), the LHS of Eq. (9a) and Eq. (9d) are always real, Eq. (9b) and Eq. (9c) are conjugate complex values, then we rewrite the state vector in a simplified format as
@@
with the dimension $23N-$ 2, where $\mathbb{X}_{n}(t)=\left[X_{n}^{R}[1],X_{n}^{R}[2],X_{n}^{I}[2]\right]$, $X_{n}^{R}$ and $X_{n}^{I}$, $a^{R}$ and $a^{I}$ represent the real and imaginary parts of $X_{n}(t)$ and $\langle a^{{\dagger}}a\rangle$ in Eq. (11) respectively, $X_{n}^{R}[1]$ represents the first element of the vector $X_{n}^{R}$, and similar for other elements in the vector $\mathbb{X}(t)$. $X_{n}^{I}[1]=a^{I}=0$, $X_{n}^{R}[3]=X_{n}^{R}[2]$ and $X_{n}^{I}[3]=-X_{n}^{I}[2]$. Then the evolution of $\mathbb{X}(t)$ reads
@@
where we take $F_{n}(t)\kappa_{n}\triangleq\mathcal{R}_{n}(t)+i\mathcal{I}_{n}(t)$, $F_{n}^{*}(t)\kappa_{n}\triangleq\mathcal{R}_{n}(t)-i\mathcal{I}_{n}(t)$ for simplification.
@@
$\mathbb{A}_{n}(t)=\begin{bmatrix}-2\bar{\mathcal{R}}_{n}&2\hat{g}_{n}&2\breve{g}_{n}\\ -\hat{g}_{n}&-\bar{\mathcal{R}}_{n}&\bar{\mathcal{I}}_{n}-\Delta_{n}\\ -\breve{g}_{n}&\Delta_{n}-\bar{\mathcal{I}}_{n}&-\bar{\mathcal{R}}_{n}\\ \end{bmatrix}$ (43)
$+\begin{bmatrix}2\bar{\mathcal{R}}_{n}-2\mathcal{R}_{n}(t)&0&0\\ 0&\bar{\mathcal{R}}_{n}-\mathcal{R}_{n}(t)&\mathcal{I}_{n}(t)-\bar{\mathcal{I}}_{n}\\ 0&\bar{\mathcal{I}}_{n}-\mathcal{I}_{n}(t)&\bar{\mathcal{R}}_{n}-\mathcal{R}_{n}(t)\\ \end{bmatrix}$
@@ III-B With external drives
When there is a drive field with the amplitude $\mathcal{E}$ applied upon the cavity, the Hamiltonian $H_{t}$ in Eq. (2) should be replaced by
$H_{t}^{\prime}$ $=H_{t}-i\mathcal{E}\left(a-a^{{\dagger}}\right).$ (44)
We denote $R_{a}=\left\langle a+a^{{\dagger}}\right\rangle$, and $I_{a}=i\left\langle a-a^{{\dagger}}\right\rangle$ with real values, and $s_{n}^{R}$ and $s_{n}^{I}$ represent the real and imaginary parts of $\left\langle\sigma^{-}_{n}\right\rangle$, respectively. Then generalized from the state vector $\mathbb{X}(t)$ in Eq. (39), we define another real-valued vector $\tilde{\mathbb{X}}$ as
@@
$\mathbb{B}(t)=\begin{bmatrix}\mathbb{B}_{1}(t)&0&\cdots&0&\tilde{\mathbb{D}}_{1}(t)\\ 0&\mathbb{B}_{2}(t)&\cdots&0&\tilde{\mathbb{D}}_{2}(t)\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\mathbb{B}_{N-1}(t)&\tilde{\mathbb{D}}_{N-1}(t)\\ \tilde{\mathbb{C}}_{1}(t)&\tilde{\mathbb{C}}_{2}(t)&\cdots&\tilde{\mathbb{C}}_{N-1}(t)&\tilde{\mathbb{P}}\end{bmatrix},$ (46)
$\mathbb{B}_{n}=$ (47)
$\begin{bmatrix}-2\mathcal{R}_{n}&2\hat{g}_{n}&2\breve{g}_{n}&0&0\\ -\hat{g}_{n}&-\mathcal{R}_{n}&\mathcal{I}_{n}-\Delta_{n}&\mathcal{E}&0\\ -\breve{g}_{n}&\Delta_{n}-\mathcal{I}_{n}&-\mathcal{R}_{n}&0&-\mathcal{E}\\ 0&0&0&-\mathcal{R}_{n}&\mathcal{I}_{n}\\ 0&0&0&-\mathcal{I}_{n}&-\mathcal{R}_{n}\end{bmatrix},$
~$\tilde{\mathbb{C}}_{n}$ $=\begin{bmatrix}0&-2\hat{g}_{n}&-2\breve{g}_{n}&0&0\\ 0&0&0&-2\hat{g}_{n}&2\breve{g}_{n}\\ 0&0&0&2\breve{g}_{n}&2\hat{g}_{n}\end{bmatrix},$ (48)
@@
Figure 4: The non-Markovian dynamics of quantum system with one two-level atom in the cavity with applied drives.
We take the two-level atom as an example to illustrate how the dynamics can be influenced by the non-Markovian environment, coupling strength between the atom and cavity, and the external drive applied on the cavity, which are compared in Fig. 4. In all the simulations, $\omega_{1}=37.7$GHz, $\kappa_{1}=0.31$GHz, the time step for simulations and the non-Markovian environmental parameter settings are the same as in Fig. 3. In Fig. 4(a), the cavity is resonant with the atom, and we compared how the dynamics is influenced by the coupling strength $g_{1}$ when $\mathcal{E}=10$MHz. It can be seen that when the cavity is driven, there can be multi-photon states in the cavity when the coupling strength between the atom and cavity is weak. In Fig. 4(b), the atom is initially in a superposition state, $g_{1}=20$MHz, $\Delta_{1}=-20$MHz, there can also be multi-photon states when the cavity is driven.
@@ IV Quantum measurement feedback control in the non-Markovian cavity-QED system
When the quantum system is measured via the cavity output as in Fig. 1, the feedback control can be designed according to the measurement result [60]
$I_{c}(t)$ $=\sqrt{2}\langle x\rangle+\xi(t)/\sqrt{\eta},$ (51)
@@
$H_{\rm fb}(t)$ $=GI_{c}(t)=G\left(\sqrt{2}\langle x\rangle+\frac{1}{\sqrt{\eta}}\xi(t)\right),$ (52)
where $G=G^{{\dagger}}$ is the feedback operator. We assume that the detection efficiency of the measurement apparatus is ideal as $\eta=1$.
@@
$\mathrm{d}\rho=-i\left[\bar{H}+H_{D},\rho\right]\mathrm{d}t+\sum_{n=1}^{N-1}\left[\left(F_{n}+F_{n}^{*}\right)L_{n}\rho L_{n}^{{\dagger}}\right.$ (53)
$\left.-\left(F_{n}L_{n}^{{\dagger}}L_{n}\rho+F_{n}^{*}\rho L_{n}^{{\dagger}}L_{n}\right)\right]\mathrm{d}t+\kappa\left[F_{a}+F_{a}^{*}\right]a\rho a^{{\dagger}}\mathrm{d}t$
$-\kappa\left[F_{a}a^{{\dagger}}a\rho+F_{a}^{*}\rho a^{{\dagger}}a\right]\mathrm{d}t-ig_{f}\left[G,a\rho+\rho a^{{\dagger}}\right]\mathrm{d}t$
$-\frac{g_{f}^{2}}{2}\left[G,\left[G,\rho\right]\right]\mathrm{d}t+\mathcal{H}[a]\rho\mathrm{d}W-ig_{f}\left[G,\rho\right]\mathrm{d}W,$
@@ IV-A Feedback control with cavity operators
In the following, we assume that the interaction between the cavity and environment is Markovian, by taking $F_{a}\equiv\kappa/2$. Additionally, the feedback operator can be designed as [63]
@@
where the position operator $x$ is given by Eq. (51) and the momentum operator $p=i\left(a^{{\dagger}}-a\right)/\sqrt{2}$ with $xp-px=i$. Besides, Eq. (54) is in the rotating frame with respect to the feedback driving frequency $\omega_{d}$ from the original format $\tilde{G}=\frac{\beta_{x}-i\beta_{p}}{\sqrt{2}}ae^{i\omega_{d}t}+\frac{\beta_{x}+i\beta_{p}}{\sqrt{2}}a^{{\dagger}}e^{-i\omega_{d}t}$ [64]. Denote $\Delta=\omega_{c}-\omega_{d}$, then the dynamics of the mean-value of the operators $x$ and $p$ can be derived based on Eq. (53) without $H_{D}$ as
$\langle\dot{x}\rangle=\Delta\langle p\rangle-\frac{\kappa}{2}\langle x\rangle$
$~{}~{}~{}~{}~{}~{}~{}-i\sum_{n}\frac{g_{n}e^{-i\delta t}}{\sqrt{2}}\left\langle\sigma^{-}_{n}\right\rangle+i\sum_{n}\frac{g_{n}e^{i\delta t}}{\sqrt{2}}\left\langle\sigma^{+}_{n}\right\rangle$
$~{}~{}~{}~{}~{}~{}~{}+\sqrt{2}g_{f}\beta_{p}\langle x\rangle+\sqrt{2}\left(V_{x}-\frac{1}{2}\right)\xi(t)+g_{f}\beta_{p}\xi(t),$ (55a)
$\langle\dot{p}\rangle=-\Delta\langle x\rangle-\frac{\kappa}{2}\langle p\rangle$
$~{}~{}~{}~{}~{}~{}~{}-\sum_{n}\frac{g_{n}e^{-i\delta t}}{\sqrt{2}}\left\langle\sigma^{-}_{n}\right\rangle-\sum_{n}\frac{g_{n}e^{i\delta t}}{\sqrt{2}}\left\langle\sigma^{+}_{n}\right\rangle$
$~{}~{}~{}~{}~{}~{}~{}-\sqrt{2}g_{f}\beta_{x}\langle x\rangle+\sqrt{2}V_{xp}\xi(t)-g_{f}\beta_{x}\xi(t),$ (55b)
@@
$\langle\dot{a}\rangle=-i\Delta\langle a\rangle-i\sum_{n}g_{n}e^{-i\delta t}\langle\sigma^{-}_{n}\rangle-\frac{\kappa}{2}\langle a\rangle$
$~{}~{}~{}~{}~{}~{}~{}-ig_{f}\left(\beta_{x}+i\beta_{p}\right)\langle x\rangle-\frac{ig_{f}}{\sqrt{2}}\left(\beta_{x}+i\beta_{p}\right)\xi(t)$
$~{}~{}~{}~{}~{}~{}~{}+\left[\left\langle\left(a+a^{{\dagger}}\right)a\right\rangle-\left\langle a+a^{{\dagger}}\right\rangle\left\langle a\right\rangle\right]\xi(t),$ (55d)
where $V_{x}$ is the variance of the position operator as $V_{x}=\left\langle x^{2}\right\rangle-\langle x\rangle^{2}$, $V_{xp}$ is the covariance between position and momentum as $V_{xp}=\left(\langle xp\rangle+\langle px\rangle\right)/2-\langle x\rangle\langle p\rangle$ [63], and $F_{n}(t)$ is governed by Eq. (9e).
Then based on the noise components in Eqs. (55a) and (55b), we can derive the following proposition on the relationship between feedback control and the effect of noise in quantum dynamics.
@@
The proposition can be easily derived by the noise components in Eqs. (55a) and (55b), thus further details are omitted. ∎
@@
$\dot{V}_{x}=2\Delta V_{xp}-\kappa V_{x}^{\prime}+2\sqrt{2}g_{f}\beta_{p}V_{x}^{\prime}+g_{f}^{2}\beta_{p}^{2}$
$~{}~{}~{}~{}~{}~{}~{}-\left(\sqrt{2}V_{x}^{\prime}+g_{f}\beta_{p}\right)^{2},$ (56a)
$\dot{V}_{xp}=\Delta V_{p}-\Delta V_{x}-\kappa V_{xp}+\sqrt{2}g_{f}\beta_{p}V_{xp}$
$~{}~{}~{}~{}~{}~{}~{}~{}-\sqrt{2}g_{f}\beta_{x}V_{x}+\frac{\sqrt{2}}{2}g_{f}\beta_{x}-g_{f}^{2}\beta_{x}\beta_{p}$
$~{}~{}~{}~{}~{}~{}~{}~{}-\left(\sqrt{2}V_{xp}-g_{f}\beta_{x}\right)\left(\sqrt{2}V_{x}^{\prime}+g_{f}\beta_{p}\right),$ (56b)
$\dot{V}_{p}=-2\Delta V_{xp}-\kappa V_{P}+\frac{\kappa}{2}-2\sqrt{2}g_{f}\beta_{x}V_{xp}+g_{f}^{2}\beta_{x}^{2}$
$~{}~{}~{}~{}~{}~{}~{}-\left(\sqrt{2}V_{xp}-g_{f}\beta_{x}\right)^{2},$ (56c)

### v4

@@ III Time-varying linear quantum control dynamics
In the following, we only consider the circumstance that $4\kappa_{n}S_{n}-Q_{n}^{2}\leq 0$, thus the interaction between the atom and environment finally converges to be Markovian. Return to Eq. (11), we analyze how the quantum dynamics can be influenced by detunings between the atom and cavity, and drive fields applied upon the quantum system.
@@
$=$ $\begin{bmatrix}-L_{n}^{(1)}&-i\tilde{g}_{n}&i\tilde{g}_{n}\\ -i\tilde{g}_{n}&-L_{n}^{(2)}&0\\ i\tilde{g}_{n}&0&-L_{n}^{(3)}\end{bmatrix}+{\rm{diag}}\left(L_{n}^{(1)}-\left(F_{n}+F_{n}^{*}\right),L_{n}^{(2)}-F_{n}^{*},L_{n}^{(3)}-F_{n}\right),$
@@
For the special case with $g_{n}=0$, where the multi-level atom only interacts with the non-Markovian environment with $C_{n}=D_{n}^{\rm T}=\textbf{0}_{1\times 3}$, we can construct the vector representing the populations of a multi-level atom as
$\tilde{X}(t)=\left[\langle\sigma^{+}_{1}\sigma^{-}_{1}\rangle,\cdots,\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle,\cdots,\langle\sigma^{+}_{N-1}\sigma^{-}_{N-1}\rangle\right]^{\rm T}.$
$\tilde{X}(t)$ is governed by the following real-valued equation if initially $\tilde{X}(0)$ is real,
$\frac{\mathrm{d}}{\mathrm{d}t}\tilde{X}(t)=\left(\textbf{L}+\textbf{P}(t)\right)\tilde{X}(t),$ (34)
@@
which is similar to the quantum dynamics with a time-varying control input [60], then we can derive the following results.
@@
The convergence of quantum dynamics to a Markovian regime means the convergence of $F_{n}(t)$ according to Remark 1. Consequently, the time-varying components in Eq. (34) converge to zero. In this case, $\lim_{t\rightarrow\infty}\textbf{P}(t)=\textbf{0}_{(N-1)\times(N-1)}$. Thus $\Phi\left(t,t_{0}\right)$ is bounded because $\lim_{t\rightarrow\infty}\partial_{t}\Phi\left(t,t_{0}\right)=0$. ∎
@@
where the non-Markovian interaction between the atom and the environment can influence the integral kernel in Eq. (36), and further influence the convergence rate and steady value of $\tilde{X}$. The dynamics of the equation with the format of Eq. (36) has been introduced in Ref. [61]. Generalized from Theorem 2 in Ref. [61] (Chapter 2, Page 36), we have the following corollary.
@@ III-A $\Delta_{n}\neq 0$, $g_{n}\neq 0$
In the following, we consider the most general case by separating $X(t)$ into its real and imaginary parts. Considering that in Eq. (9), $\langle\sigma^{+}_{n}\sigma^{-}_{n}\rangle$ and $\langle a^{{\dagger}}a\rangle$ are both real, Eq. (9b) and Eq. (9c) are conjugate complex values, then we rewrite the state vector in a simplified format as
@@
with the dimension 3N23N-2, where $\mathbb{X}_{n}(t)=\left[X_{n}^{R}[1],X_{n}^{R}[2],X_{n}^{I}[2]\right]$, $X_{n}^{R}$ and $X_{n}^{I}$, $a^{R}$ and $a^{I}$ represent the real and imaginary parts of $X_{n}(t)$ and $\langle a^{{\dagger}}a\rangle$ in Eq. (10) respectively, $X_{n}^{R}[1]$ represents the first element of the vector $X_{n}^{R}$, and similar for other elements in the vector $\mathbb{X}(t)$. $X_{n}^{I}[1]=a^{I}=0$, $X_{n}^{R}[3]=X_{n}^{R}[2]$ and $X_{n}^{I}[3]=-X_{n}^{I}[2]$. Then the evolution of $\mathbb{X}(t)$ reads
@@
where we take $F_{n}(t)\kappa_{n}\triangleq\mathcal{R}_{n}(t)+i\mathcal{I}_{n}(t)$ for simplification.
@@
$\mathbb{A}_{n}(t)$ $=\begin{bmatrix}-2\bar{\mathcal{R}}_{n}&2\hat{g}_{n}&2\breve{g}_{n}\\ -\hat{g}_{n}&-\bar{\mathcal{R}}_{n}&\bar{\mathcal{I}}_{n}-\Delta_{n}\\ -\breve{g}_{n}&\Delta_{n}-\bar{\mathcal{I}}_{n}&-\bar{\mathcal{R}}_{n}\\ \end{bmatrix}+\begin{bmatrix}2\bar{\mathcal{R}}_{n}-2\mathcal{R}_{n}(t)&0&0\\ 0&\bar{\mathcal{R}}_{n}-\mathcal{R}_{n}(t)&\mathcal{I}_{n}(t)-\bar{\mathcal{I}}_{n}\\ 0&\bar{\mathcal{I}}_{n}-\mathcal{I}_{n}(t)&\bar{\mathcal{R}}_{n}-\mathcal{R}_{n}(t)\\ \end{bmatrix}$ (41)
@@ III-B With external drives
When there is a drive field with the amplitude $\mathcal{E}$ applied upon the cavity, the Hamiltonian $\bar{H}_{I}$ in Eq. (2) should be replaced by
$H^{\prime}$ $=\bar{H}_{I}-i\mathcal{E}\left(a-a^{{\dagger}}\right).$ (42)
We denote $R_{a}=\left\langle a+a^{{\dagger}}\right\rangle$, $I_{a}=i\left\langle a-a^{{\dagger}}\right\rangle$ with real values, $s_{n}^{R}$ and $s_{n}^{I}$ represent the real and imaginary parts of $\left\langle\sigma^{-}_{n}\right\rangle$, respectively. Then generalized from the state vector $\mathbb{X}(t)$ in Eq. (37), we define another real-valued vector $\tilde{\mathbb{X}}$ as
@@
$\mathbb{B}(t)=\begin{bmatrix}\mathbb{B}_{1}(t)&0&\cdots&0&\tilde{\mathbb{D}}_{1}(t)\\ 0&\mathbb{B}_{2}(t)&\cdots&0&\tilde{\mathbb{D}}_{2}(t)\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&\mathbb{B}_{N-1}(t)&\tilde{\mathbb{D}}_{N-1}(t)\\ \tilde{\mathbb{C}}_{1}(t)&\tilde{\mathbb{C}}_{2}(t)&\cdots&\tilde{\mathbb{C}}_{N-1}(t)&\tilde{\mathbb{P}}\end{bmatrix}$, ($4$)
$\mathbb{B}_{n}=\begin{bmatrix}-2\mathcal{R}_{n}&2\hat{g}_{n}&2\breve{g}_{n}&0&0\\ -\hat{g}_{n}&-\mathcal{R}_{n}&\mathcal{I}_{n}-\Delta_{n}&\mathcal{E}&0\\ -\breve{g}_{n}&\Delta_{n}-\mathcal{I}_{n}&-\mathcal{R}_{n}&0&-\mathcal{E}\\ 0&0&0&-\mathcal{R}_{n}&\mathcal{I}_{n}\\ 0&0&0&-\mathcal{I}_{n}&-\mathcal{R}_{n}\end{bmatrix}$, (45)
$\tilde{\mathbb{C}}_{n}$ $=\begin{bmatrix}0&-2\hat{g}_{n}&-2\breve{g}_{n}&0&0\\ 0&0&0&-2\hat{g}_{n}&2\breve{g}_{n}\\ 0&0&0&2\breve{g}_{n}&2\hat{g}_{n}\end{bmatrix}$, (46)
@@
Figure 4: Non-Markovian dynamics of the quantum system with one two-level atom coupled to the cavity with applied drives.
We take the two-level atom as an example to illustrate how the dynamics can be influenced by the non-Markovian environment, coupling strength between the atom and cavity, and the external drive applied on the cavity, which are compared in Fig. 4. In all the simulations, $\omega_{1}=37.7$GHz, $\kappa_{1}=0.31$GHz, the time step for simulations and the non-Markovian environmental parameter settings are the same as in Fig. 3. In Fig. 4(a), the cavity is resonant with the atom, and we compared how the dynamics is influenced by the coupling strength $g_{1}$ when $\mathcal{E}=10$MHz. It can be seen that when the cavity is driven, there can be multi-photon states in the cavity if the coupling strength between the atom and cavity is weak. In Fig. 4(b), the atom is initially in a superposition state, $g_{1}=20$MHz, $\Delta_{1}=-20$MHz, there can also be multi-photon states when the cavity is driven.
@@ IV Quantum measurement feedback control in the non-Markovian cavity-QED system
When the quantum system is measured via the cavity output as in Fig. 1, the feedback control can be designed according to the measurement result [1, 63]
$I_{c}(t)$ $=\sqrt{2}\langle x\rangle+\frac{\xi(t)}{\sqrt{\eta}},$ (49)
@@
$H_{\rm fb}(t)$ $=I_{c}(t)G,$ (50)
where $G=G^{{\dagger}}$ is the feedback operator. We assume that the detection efficiency of the measurement apparatus is ideal as $\eta=1$, and see Ref. [64] for the case $\eta<1$ with a similar format.
@@
$\mathrm{d}\rho=$ $-i\left[\bar{H}+H_{D},\rho\right]\mathrm{d}t+\sum_{n=1}^{N-1}\left[\left(F_{n}+F_{n}^{*}\right)L_{n}\rho L_{n}^{{\dagger}}-\left(F_{n}L_{n}^{{\dagger}}L_{n}\rho+F_{n}^{*}\rho L_{n}^{{\dagger}}L_{n}\right)\right]\mathrm{d}t$ (51)
$+\kappa\left(F_{a}+F_{a}^{*}\right)a\rho a^{{\dagger}}\mathrm{d}t-\kappa\left(F_{a}a^{{\dagger}}a\rho+F_{a}^{*}\rho a^{{\dagger}}a\right)\mathrm{d}t$
$-ig_{f}\left[G,a\rho+\rho a^{{\dagger}}\right]\mathrm{d}t-\frac{g_{f}^{2}}{2}\left[G,\left[G,\rho\right]\right]\mathrm{d}t+\mathcal{H}[a]\rho\mathrm{d}W-ig_{f}\left[G,\rho\right]\mathrm{d}W,$
@@ IV-A Feedback control with cavity operators
In this subsection, we assume that the interaction between the cavity and environment is Markovian with the decay rate $\kappa$. The feedback operator can be designed as [67]
@@
where the position operator $x$ is given by Eq. (49) and the momentum operator $p=i\left(a^{{\dagger}}-a\right)/\sqrt{2}$ with xppx=ixp-px=i. Besides, Eq. (52) is in the rotating frame with respect to the feedback driving frequency $\omega_{d}$ from the original format $\tilde{G}=\left[\left(\beta_{x}-i\beta_{p}\right)ae^{i\omega_{d}t}+\left(\beta_{x}+i\beta_{p}\right)a^{{\dagger}}e^{-i\omega_{d}t}\right]/\sqrt{2}$ [68]. Denote $\Delta=\omega_{c}-\omega_{d}$, then the dynamics of the mean-value of the operators $x$ and $p$ can be derived based on Eq. (51) without $H_{D}$ as
$\langle\dot{x}\rangle=\Delta\langle p\rangle-\frac{\kappa}{2}\langle x\rangle-i\sum_{n}\frac{g_{n}e^{-i\delta t}}{\sqrt{2}}\left\langle\sigma^{-}_{n}\right\rangle+i\sum_{n}\frac{g_{n}e^{i\delta t}}{\sqrt{2}}\left\langle\sigma^{+}_{n}\right\rangle+\sqrt{2}g_{f}\beta_{p}\langle x\rangle+\sqrt{2}\left(V_{x}-\frac{1}{2}\right)\xi(t)+g_{f}\beta_{p}\xi(t),$ (53a)
$\langle\dot{p}\rangle=-\Delta\langle x\rangle-\frac{\kappa}{2}\langle p\rangle-\sum_{n}\frac{g_{n}e^{-i\delta t}}{\sqrt{2}}\left\langle\sigma^{-}_{n}\right\rangle-\sum_{n}\frac{g_{n}e^{i\delta t}}{\sqrt{2}}\left\langle\sigma^{+}_{n}\right\rangle-\sqrt{2}g_{f}\beta_{x}\langle x\rangle+\sqrt{2}V_{xp}\xi(t)-g_{f}\beta_{x}\xi(t),$ (53b)
@@
$\langle\dot{a}\rangle=-i\Delta\langle a\rangle-i\sum_{n}g_{n}e^{-i\delta t}\langle\sigma^{-}_{n}\rangle-\frac{\kappa}{2}\langle a\rangle-ig_{f}\left(\beta_{x}+i\beta_{p}\right)\langle x\rangle$
$~~~~~~~-ig_{f}\left(\beta_{x}+i\beta_{p}\right)\xi(t)/\sqrt{2}+\left[\left\langle\left(a+a^{{\dagger}}\right)a\right\rangle-\left\langle a+a^{{\dagger}}\right\rangle\left\langle a\right\rangle\right]\xi(t),$ (53d)
where $V_{x}=\left\langle x^{2}\right\rangle-\langle x\rangle^{2}$ is the variance of the position operator, $V_{p}$ is that for the momentum operator, and $V_{xp}=\left(\langle xp\rangle+\langle px\rangle\right)/2-\langle x\rangle\langle p\rangle$ is the covariance between position and momentum [67]. Then based on the stochastic components in Eqs. (53a) and (53b), we have the following proposition.
@@
The proposition can be easily derived by the noise components in Eqs. (53a) and (53b), thus more details are omitted. ∎
@@
$\dot{V}_{x}=2\Delta V_{xp}-\kappa V_{x}^{\prime}+2\sqrt{2}g_{f}\beta_{p}V_{x}^{\prime}+g_{f}^{2}\beta_{p}^{2}-\left(\sqrt{2}V_{x}^{\prime}+g_{f}\beta_{p}\right)^{2},$ (54a)
$\dot{V}_{xp}=\Delta V_{p}-\Delta V_{x}-\kappa V_{xp}+\sqrt{2}g_{f}\beta_{p}V_{xp}-\sqrt{2}g_{f}\beta_{x}V_{x}+\frac{\sqrt{2}}{2}g_{f}\beta_{x}-g_{f}^{2}\beta_{x}\beta_{p}-\left(\sqrt{2}V_{xp}-g_{f}\beta_{x}\right)\left(\sqrt{2}V_{x}^{\prime}+g_{f}\beta_{p}\right),$ (54b)
$\dot{V}_{p}=-2\Delta V_{xp}-\kappa V_{P}+\frac{\kappa}{2}-2\sqrt{2}g_{f}\beta_{x}V_{xp}+g_{f}^{2}\beta_{x}^{2}-\left(\sqrt{2}V_{xp}-g_{f}\beta_{x}\right)^{2},\$(54c)